OpenStudy (anonymous):

is friction between the block and the table. The suspended 3 kg mass on the left is moving up, the 2 kg mass slides to the right on the table, and the suspended mass 6 kg on the right is moving down. The acceleration of gravity is 9 . 8 m /

OpenStudy (fifciol):

 Newton's second law: \[T_1-m_1g=m_1a\]\[T_2-T_1-F_r=m_2a\]\[m_3g-T_2=m_3a\] Solve for Fr